# CORE-BEHRT A Carefully Optimized and Rigorously Evaluated BEHRT

### v2

@@ Abstract
BERT-based models for Electronic Health Records (EHR) have surged in popularity following the release of BEHRT and Med-BERT. Subsequent models have largely built on these foundations despite the fundamental design choices of these pioneering models remaining underexplored. To address this issue, we introduce CORE-BEHRT, a Carefully Optimized and Rigorously Evaluated BEHRT. Through incremental optimization, we isolate the sources of improvement for key design choices, giving us insights into the effect of data representation and individual technical components on performance. Evaluating this across a set of generic tasks (death, pain treatment, and general infection), we showed that improving data representation can increase the average downstream performance from 0.785 to 0.797 AUROC ($p<10^{-7}$), primarily when including medication and timestamps. Improving the architecture and training protocol on top of this increased average downstream performance to 0.801 AUROC ($p<10^{-7}$). We then demonstrated the consistency of our optimization through a rigorous evaluation across 25 diverse clinical prediction tasks. We observed significant performance increases in 17 out of 25 tasks and improvements in 24 tasks, highlighting the generalizability of our findings. Our findings provide a strong foundation for future work and aim to increase the trustworthiness of BERT-based EHR models.
Keywords BEHRT, Med-BERT, BERT, Transformers, EHR, Risk Prediction, Deep Learning, Real World Data
1 Introduction
@@
In recent years, leveraging Electronic Health Records (EHR) to predict patient outcomes has gained significant interest. Initially, progress was gradual due to a select number of available machine learning methods, such as logistic regression and random forest (MJ et al., 1997; Khalilia et al., 2011). However, the advent of deep learning, coupled with a drastic increase in available EHR data (Slawomirski et al., 2023), marked a pivotal shift and enabled the extraction of more complex patterns and insights.
@@
A notable development in utilizing EHR data more effectively has been the adoption of architectures originally developed for natural language processing (NLP) and computer vision. This has been exemplified by the use of recurrent (Choi et al., 2015, 2016a, 2016b) and convolutional (Nguyen et al., 2016) neural networks on EHR data. The introduction of the transformer architecture (Vaswani et al., 2017), known for its versatility across different data types, significantly impacted EHR analysis. Specifically, building on the seminal work of BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018), the field has seen the emergence of transformer-based models tailored to EHR data (Li et al., 2020; Rasmy et al., 2021). These early models inspired a wave of subsequent research introducing a variety of modifications touching on input data representation, model architecture, downstream tasks, and evaluation metrics as seen in works like BRLTM (Y et al., 2021) and Hi-BEHRT (Li et al., 2022) or use ICU-specific data (Zhang et al., 2022). This heterogeneity complicates direct comparisons and makes it untractable to pinpoint the sources of improvement.
@@
Figure 1: This figure depicts the selection process of the best configuration. Non-selected configurations are depicted in a lighter shade, indicating they were considered but not chosen based on their impact on the average Area Under the Receiver Operating Characteristic (AUROC), across three prediction tasks, whose prediction windows are denoted in parentheses. Models were trained in five-fold cross-validation and evaluated on one test set per task. An extension of these experiments for masking ratios and pooling strategies is demonstrated in Appendix I.
@@
Despite the parallels with NLP, it is also important to acknowledge the distinct characteristics of EHR data, notably its strong temporal aspect, irregular time intervals, and the distinct distribution of sequence lengths (Rasmy et al., 2021). These disparities necessitate a comprehensive evaluation of components that have shown success in NLP within the context of EHR. Furthermore, design choices take a weighted importance in healthcare compared to NLP, as even minor performance improvements can significantly impact patient outcomes.
In contrast to fields like radiology, where deep learning models increasingly find their way into clinical practice (Yasaka and Abe, 2018; Montagnon et al., 2020; Lauritzen et al., 2023), there is a lack of trust towards EHR-based models, a cornerstone for clinical adoption (Asan et al., 2020). Therefore, by providing insights into these models, we increase their trustworthiness and make a step towards the integration of BERT-based EHR models into clinical workflows.
@@
Main Contributions
Weak and outdated foundations, coupled with the heterogeneity in newer models, pose challenges in making direct comparisons of BERT-based EHR models. This creates a lack of trust, a major hurdle to clinical adoption, and restricts these models’ use to research settings. We address this by conducting a careful optimization rigorously evaluated across a diverse set of tasks to increase insights into and confidence towards BERT-based EHR models. Our main contributions consist of:
  • Key insights in data representation, model architecture, and training protocol through careful optimization, robustly showing model improvements
  • A comprehensive and rigorous evaluation across 25 diverse clinical prediction tasks showcasing the consistency of the optimized model
@@ 2.2 Transformers in NLP
3 Cohort
@@
The data consists of EHRs from 1,825 million patients, prospectively recorded during the period from 2016 to 2022 in the Capital Region of Denmark. We consider only patients with at least one code eligible, which results in 1,805 million patients. For FT, we deem at least one code to be a proper input for binary classification but use patients with $\geq 3$ codes for PT to ensure a proper input for the MLM objective task. An overview of the selection pipeline can be seen in Appendix F.
@@ 3.1 Cohort Selection
To test generalizability across a wide range of tasks while managing computational demands, we maintained a manageable cohort size for the main experiments. However, we also performed "maximizing performance" experiments, where we extended the cohorts, including PT patients in the FT set for select tasks. Details of these tasks and the expanded cohort sizes are provided in Table 3. For model selection, we require each patient in the PT cohort to have at least three diagnoses to ensure a fair comparison between the diagnosis only and the diagnosis+medication models. For the final evaluation, we require at least three of any (diagnosis or medication) codes, to accurately reflect the increased cohort when including medication.
For each condition, we defined a relevant age and/or sex and a prediction window that we use for censoring, given in Table 1 for the optimization experiments and in Table 4 for the generalization assessment. The full age distributions for the cohorts before age group selection are given in Appendix H. We consider patients to be positive when having at least one of the codes given in Appendix K and negative otherwise. We address imbalance and potential information leakage using the methods described in subsection 4.4.
3.2 Data Extraction
We extract all medical events, with an event consisting of three main variables: concept, timestamp, and admission ID. These events undergo minimal data processing through a simple three-stage process of replacing missing ICD10 codes with textual descriptions, attempting to impute missing values, only when confident, before excluding any events with missing info or ages outside a normal range. We leave a more detailed description in Appendix C.
3.3 Feature Choices
@@
4 Methods
Figure 2: Overview of the iterative optimization procedure, highlighting adapted settings in bold. We augmented the input data with medication codes, patient sex, event timestamps, full International Classification of Diseases (ICD-10), and Anatomical Therapeutic Chemical (ATC) codes. Improvements to the model include integrating time2vec embeddings, Rotary Position Embeddings (RoPE), and the SwiGLU activation function. For pre-training (PT), we investigated various masking ratios and Med-BERT’s secondary PT task: predicting prolonged length of stay in the hospital. We investigated various pooling strategies for fine-tuning (FT), finalizing on Bidirectional Gated Recurrent Units (BiGRU) based on its superior results.
We will introduce our base model and the modifications tested, separated into a data presentation and a technical part. The data representation part investigates the impact of representing commonly available background EHR data with little to no implementation overhead. The technical part investigates the impact of improved architecture, embeddings, and training schemes, primarily relevant to machine learning practitioners. Modifications are introduced incrementally, evaluating their impact on the performance of three generic downstream tasks and including them when there is an increase in average Area Under the Receiver Operating Characteristic (AUROC). An overview of this procedure is shown in Figure 2. Next, we describe the optimization procedure before defining the various downstream tasks and close with the hyperparameters of our architecture and optimization.
4.1 Base Model
We used a slightly adapted BEHRT as our starting point. BEHRT’s input data consist of medical events with CLS (classification) and SEP (separator) tokens, age in years, and visits denoted by "positions" and binary visit encodings (segments).
We adopted Med-BERT’s leaner configuration for hyperparameters to maintain manageable model dimensions, as BEHRT could use more layers and heads due to its smaller vocabulary size. Further details are provided in subsection 4.6.
4.2 Data Representation
We investigate different components of data representation which are easily integrated into the model and commonly ignored.
Medication Codes
We adapted the BEHRT model to include medication codes, recognizing their potential to improve predictive modeling (Wei et al., 2016). Although the inclusion of medication codes, due to their similar coding system to diagnoses, is not unconventional, it has yet to become standard practice. We consider medication a vital addition to the model, as diagnosis codes are often incomplete or inadequate (Horsky et al., 2017). For simplicity and computational feasibility, our current model includes only the medication codes, akin to how we handle diagnoses, without considering dosage information. Although our dataset also contains lab tests, vital signs, and procedures, with significantly higher event counts (672.2 million, 605.9 million, and 300.3 million events, respectively), we focus solely on the 227.6 million medication events due to their direct relevance to treatment decisions.
Full-Depth Codes and Background Sentence
We propose to include more fine-grained data by using the full-depth medical and diagnostic codes and include the patient’s sex. It is common for EHR models to limit the input to higher-level codes, such as in BEHRT (Li et al., 2020). This can be beneficial as high-level representation decreases the vocabulary size which in turn decreases model size. It may also reduce noise to represent information on a coarse level only. Despite these advantages, it can also remove critical information and might especially be detrimental when rare conditions are of interest. BEHRT mapped their codes to 301 Caliber categories, far less than the number of ICD10 codes. When utilizing level four codes, our model works with a vocabulary size of 2,475, whereas using the full codes expands the vocabulary to 17,469. This highlights the balance between model complexity and the granularity of information captured.
To include the patient’s sex in the model, we prepend a token representing male or female to their sequence. We foresee that other relevant demographic variables could be included as well.
Timestamps
Separator Tokens
4.3 Technical Modifications
Here, we introduce the technical components, such as new embeddings and architectural changes, and pre-training (PT) objectives that were tested.
Simplified Segments
Time2Vec
We suggest substituting BEHRT’s conventional age embedding with a Time2Vec (Kazemi et al., 2019) embedding. BEHRT’s current model employs embeddings that rely on discrete age in years, which can be problematic, especially in pediatric cases where changes occur on shorter timescales. These embeddings may also struggle to discern the relationship between different ages. Time2Vec models this relationship explicitly, using differently parametrized periodic functions, by giving similar representations to similar ages. Adequately understanding the relationships between ages necessitates exposure to a wide range of examples for each age. To overcome these limitations and to accommodate the timestamps mentioned previously, we propose the adoption of Time2Vec. We leave the details of Time2vec and our implementation of it in Appendix A.
Improved Transformer Recipe
Since the original transformer paper, numerous refinements have been suggested. In this study, we incorporate two such advancements from the improved transformer recipe, dubbed transformer++, that are particularly effective in boosting performance: Rotary Position Embedding (RoPE) (Su et al., 2023) and the SwiGLU activation function (Shazeer, 2020).
$\textrm{SwiGLU}(x)$ $=(\textrm{Swish}(xW_{1})\otimes xW_{2})W_{3}\,,$ (1)
$\textrm{Swish}(x)$ $=x*\mathrm{sigmoid}(x)\,.$ (2)
Pre-training Objective
The training scheme of BEHRT plays a major role in the performance, as the strength of these models largely comes from PT, specifically Masked Language Modelling (MLM). Here, BEHRT copied the setup from BERT, masking 15% of the tokens—referred to as the masking ratio (MR)—using the 80-10-10 strategy (80% mask, 10% random word, 10% unchanged). We have yet to see any evidence-based reason for picking 15% in EHR, and other modalities (images, audio, and videos) have much higher MRs (Dosovitskiy et al., 2021; He et al., 2021; Baevski et al., 2020; Tong et al., 2022). It has also been seen that larger models tend to favor larger MRs (Wettig et al., 2023).
Furthermore, we tested the second binary classification PT objective introduced by Med-BERT, in which a prediction is made on whether the patient has a prolonged length of stay in the hospital ($>7$ days) for each patient. This task was run alongside MLM.
Pooling Strategies
4.4 Optimization Procedure
For evaluation, we use three generic downstream tasks for generalizable performance: one-year death prediction for patients older than 60 y.o., one-month pain treatment prediction, and one-week infection prediction; the corresponding definitions are given in Appendix K. Death encapsulates the final outcome and can signal the overall health status, encompassing various conditions and factors affecting patient well-being. Pain treatment is very frequent and covers a wide range of causes. It can be both mild and severe, indicative of many underlying conditions, and thus quite general. Infection points to a compromised immune system, which also tells us about the general condition of a patient.
Our three tasks constitute a generalizable evaluation by covering a broad range of health conditions. To compare the performance between models, we mainly considered the area under the receiver operating curve (AUROC) but also reported the area under the precision-recall curve (AUPRC) in Appendix J. For each downstream task, we perform five-fold cross-validation (CV) evaluating on a test set and compute the mean and standard deviation. The downstream task was a binary classification and was constructed as follows. For patients with positive outcomes, we used the outcome of interest as the index date. We then excluded all events within a specified prediction window leading up to the index date, with the window tailored to each downstream task. For patients with negative outcomes, we assign random censoring dates matching the distribution of the positive cases to prevent information leakage. To account for the imbalance in the dataset we sampled with the inverse square root of the frequency of each class (Mahajan et al., 2018; Mikolov et al., 2013).
4.5 Assessing Model Generalizability
We also covered conditions with different associations to overall health and underlying factors. For instance, the strong correlation between stroke and overall health (Arboix, 2015) contrasts with the more complex interplay of genetic and environmental factors in schizophrenia (Stilo and Murray, 2019; Henriksen et al., 2017), which presents a subtler EHR data signature. We also covered conditions with a limited range of causes, such as anaphylaxis, which primarily involves drug-related causes, and those that can have a wide range of underlying conditions, such as sleep disorder. The conditions are listed in Table 4 together with prediction window and age groups, which were chosen based on relevance and considering the age distributions provided in Appendix H.
This structured approach to task selection is designed to demonstrate the generalizability of our model in identifying a wide range of health outcomes, underscoring its potential utility in diverse clinical scenarios without having to revisit the model choices for every new downstream task. We construct each downstream task as outlined in the previous subsection 4.4 and perform hypothesis testing as described in Appendix B.
4.6 Hyperparameters
To ensure a fair comparison, we standardized the hyperparameters across all experiments. We aimed to facilitate stable model convergence across all models while keeping manageable training durations. We utilized a base model architecture consisting of six layers and six attention heads with a hidden size of 192 and an intermediate size of 64. Sequences were truncated to 512 tokens, and a batch size of 256 for PT and 512 for FT was used. Learning rates were set to $10^{-3}$ for PT and $2\cdot 10^{-5}$ for FT using a linear warmup of five epochs. Models were trained for 50 epochs during optimization and 150 epochs for the generalization experiments with an early stopping of five epochs. A more detailed overview can be found in Appendix D.
5 Results on Real Data
5.1 Evaluation Approach/Study Design

### v4

# CORE-BEHRT: A Carefully Optimized and Rigorously Evaluated BEHRT

@@ Abstract
The widespread adoption of Electronic Health Records (EHR) has significantly increased the amount of available healthcare data. This has allowed models inspired by Natural Language Processing (NLP) and Computer Vision, which scale exceptionally well, to be used in EHR research. Particularly, BERT-based models have surged in popularity following the release of BEHRT and Med-BERT. Subsequent models have largely built on these foundations despite the fundamental design choices of these pioneering models remaining underexplored. Through incremental optimization, we study BERT-based EHR modeling and isolate the sources of improvement for key design choices, giving us insights into the effect of data representation, individual technical components, and training procedure. Evaluating this across a set of generic tasks (death, pain treatment, and general infection), we showed that improving data representation can increase the average downstream performance from 0.785 to 0.797 AUROC ($p<10^{-7}$), primarily when including medication and timestamps. Improving the architecture and training protocol on top of this increased average downstream performance to 0.801 AUROC ($p<10^{-7}$). We then demonstrated the consistency of our optimization through a rigorous evaluation across 25 diverse clinical prediction tasks. We observed significant performance increases in 17 out of 25 tasks and improvements in 24 tasks, highlighting the generalizability of our results. Our findings provide a strong foundation for future work and aim to increase the trustworthiness of BERT-based EHR models.
00footnotetext: The code is available on https://github.com/mikkelfo/CORE-BEHRT
Appendix A Introduction
@@
In recent years, leveraging Electronic Health Records (EHR) to predict patient outcomes has gained significant interest. Initially, progress was gradual due to a select number of available machine learning methods, such as logistic regression and random forest (MJ et al., 1997; Khalilia et al., 2011). Although models like XGBoost (Chen and Guestrin, 2016) show impressive performance on tabular data, the field has focused largely on adopting deep learning models from natural language processing (NLP) and computer vision. These models, coupled with the drastic increase in available EHR data (Slawomirski et al., 2023), allow sequential processing of the data, taking into account its temporal aspect.
@@
This has been exemplified by the use of recurrent (Choi et al., 2015, 2016a, 2016b) and convolutional (Nguyen et al., 2016) neural networks on EHR data. The introduction of the transformer architecture (Vaswani et al., 2017), known for its versatility across different data modalities and exceptional scaling, significantly impacted EHR analysis. Specifically, building on the seminal work of BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018), the field has seen the emergence of transformer-based models tailored to EHR data (Li et al., 2020; Rasmy et al., 2021). These early models inspired a wave of subsequent research introducing a variety of modifications touching on input data representation, model architecture, downstream tasks, and evaluation metrics as seen in works like BRLTM (Y et al., 2021) and Hi-BEHRT (Li et al., 2022) or use ICU-specific data (Zhang et al., 2022). This heterogeneity complicates direct comparisons and makes it untractable to pinpoint the sources of improvement.
@@
Figure 1: This figure depicts the incremental optimization of the configurations. Non-selected configurations are depicted in a lighter shade, indicating they were considered but not chosen based on their impact on the average Area Under the Receiver Operating Characteristic (AUROC), across three prediction tasks, whose prediction windows are denoted in parentheses. Models were trained in five-fold cross-validation and evaluated on one test set per task. An extension of these experiments for masking ratios and pooling strategies is demonstrated in Appendix Q.
@@
Despite the parallels with NLP, it is also important to acknowledge the distinct characteristics of EHR data, notably its strong temporal aspect, irregular time intervals, and the distinct distribution of sequence lengths (Rasmy et al., 2021). These disparities necessitate a comprehensive evaluation of the components that have shown success in NLP in the context of EHR. Furthermore, design choices take a weighted importance in healthcare compared to NLP, as even minor performance improvements can significantly impact patient outcomes.
In contrast to fields like radiology, where deep learning models are increasingly finding their way into clinical practice (Yasaka and Abe, 2018; Montagnon et al., 2020; Lauritzen et al., 2023), there is a lack of trust towards EHR-based models, a cornerstone for clinical adoption (Asan et al., 2020). Therefore, by providing insights into these models, we increase their trustworthiness and make a step towards the integration of BERT-based EHR models into clinical workflows.
@@
Generalizable Insights about Machine Learning in the Context of Healthcare
Weak and outdated foundations, coupled with the heterogeneity in newer models, pose challenges in making direct comparisons of BERT-based EHR models. This creates a lack of trust, a major hurdle to clinical adoption, and restricts these models’ use to research settings. We address these issues by conducting a careful optimization rigorously evaluated across a diverse set of tasks to increase insights into and confidence towards BERT-based EHR models. Our main contributions consist of:
  • Key insights in the data representation, model architecture, and training protocol of EHR modeling through careful optimization
  • A comprehensive and rigorous evaluation across 25 diverse clinical prediction tasks showcasing the consistency of the optimizations
@@ B.2 Transformers in NLP
Appendix C Methods
Figure 2: Overview of the iterative optimization procedure, highlighting adapted settings in bold. Symbols denote either addition (+) or removal (-) of a setting with no symbol indicating a direct change. We augmented the input data with medication codes, patient sex, event timestamps, full International Classification of Diseases (ICD-10), and Anatomical Therapeutic Chemical (ATC) codes. Improvements to the model include integrating time2vec embeddings, Rotary Position Embeddings (RoPE), and the SwiGLU activation function. For pre-training (PT), we investigated various masking ratios and Med-BERT’s secondary PT task: predicting prolonged length of stay in the hospital. We investigated various pooling strategies for fine-tuning (FT), finalizing on Bidirectional Gated Recurrent Units (BiGRU) based on its superior results.
We will introduce our base model and the modifications tested, separated into a data presentation and a technical part. We test a few representative components in different areas, covering as much ground as possible while keeping the computational demand within our budget. The data representation part investigates the impact of representing commonly available background EHR data with little to no implementation overhead. The technical part investigates the impact of improved architecture, embeddings, and training schemes, primarily relevant to machine learning practitioners. Modifications are introduced incrementally, evaluating their impact on the performance of three generic downstream tasks and including them when there is an increase in average Area Under the Receiver Operating Characteristic (AUROC). An overview of this procedure is shown in Figure 2. Next, we describe the optimization procedure before defining the various downstream tasks and close with the hyperparameters of our architecture and optimization.
C.1 Base Model
We used a slightly adapted BEHRT as our starting point. BEHRT’s input data consist of medical events with CLS (classification) and SEP (separator) tokens, age in years, and visits denoted by ”positions” and binary visit encodings (segments).
We adopted Med-BERT’s leaner configuration for hyperparameters to maintain manageable model dimensions, as BEHRT could use more layers and heads due to its smaller vocabulary size. Further details are provided in subsection C.6.
C.2 Data Representation
We investigate different key areas of data representation by including more data sources, extending existing data, and exploring visit representation. This approach provides valuable insight into the aspects of data representation that are crucial for performance gains.
Medication Codes
We adopted the BEHRT model to include medication codes, recognizing their potential to improve predictive modeling (Wei et al., 2016). Although the inclusion of medication codes, due to their similar coding system to diagnoses, is not unconventional, it has yet to become standard practice. We consider medication a vital addition to the model, as diagnosis codes are often incomplete or inadequate (Horsky et al., 2017). For simplicity and computational feasibility, our current model includes only the medication codes, akin to how we handle diagnoses, without considering dosage information. Although our dataset also contains lab tests, vital signs, and procedures, with significantly higher event counts (672.2 million, 605.9 million, and 300.3 million events, respectively), we focus solely on the 227.6 million medication events due to their direct relevance to treatment decisions.
Full-Depth Codes and Background Sentence
We propose including more fine-grained data by using the full-depth medical and diagnostic codes and including the patient’s sex. It is common for EHR models to limit the input to higher-level codes, such as in BEHRT (Li et al., 2020). This can be beneficial as high-level representation decreases the vocabulary size which in turn decreases model size. It may also reduce noise to represent information on a coarse level only. Despite these advantages, it can also remove critical information and might especially be detrimental when rare conditions are of interest. BEHRT mapped their codes to 301 Caliber categories, far less than the number of ICD10 codes. When utilizing level four codes, our model works with a vocabulary size of 2,475, whereas using the full codes expands the vocabulary to 17,469. This highlights the balance between model complexity and the granularity of information captured.
To include the patient’s sex in the model, we prepend a token representing male or female to their sequence. We foresee that other relevant demographic variables could also be included to form a ’background sentence’.
Separator Tokens
Simplified Segments
Timestamps
C.3 Technical Modifications
Here, we introduce the technical components selected from three key technical areas: model architecture, pre-training, and finetuning strategies. The components have been chosen due to their presence and success in the EHR or NLP literature.
Time2Vec
Like (Pang et al., 2021), we suggest substituting BEHRT’s conventional age embedding with a Time2Vec (Kazemi et al., 2019) embedding. BEHRT’s current model employs embeddings that rely on discrete age in years, which can be problematic, especially in pediatric cases where changes occur on shorter timescales. These embeddings may also struggle to discern the relationship between different ages. Time2Vec models this relationship explicitly, using differently parameterized periodic functions, by giving similar representations to similar ages. Adequately understanding the relationships between ages necessitates exposure to a wide range of examples for each age. To overcome these limitations and to accommodate the timestamps mentioned previously, we propose the adoption of Time2Vec. We leave the details of Time2vec and our implementation of it in Appendix G.
Improved Transformer Recipe
Numerous refinements have been suggested since the original transformer paper, and we incorporate two such advancements from the improved transformer recipe, dubbed transformer++, that are particularly effective in boosting performance: Rotary Position Embedding (RoPE) (Su et al., 2023) and the SwiGLU activation function (Shazeer, 2020).
$\textrm{SwiGLU}(x)$ $=(\textrm{Swish}(xW_{1})\otimes xW_{2})W_{3}\,,$ (1)
$\textrm{Swish}(x)$ $=x*\mathrm{sigmoid}(x)$ (2)
Pre-training Objective
The training scheme of BEHRT plays a major role in the performance, as the strength of these models comes largely from PT, specifically Masked Language Modelling (MLM). Here, BEHRT copied the setup from BERT, masking 15% of the tokens — referred to as the masking ratio (MR) —using the 80-10-10 strategy (80% mask, 10% random word, 10% unchanged). We have yet to see any evidence-based reason for picking 15% in EHR, and other modalities (images, audio, and videos) have much higher MRs (Dosovitskiy et al., 2021; He et al., 2021; Baevski et al., 2020; Tong et al., 2022). It has also been seen that larger models tend to favor larger MRs (Wettig et al., 2023).
Furthermore, we tested the second binary classification PT objective introduced by Med-BERT (Rasmy et al., 2021), in which a prediction is made on whether the patient has a prolonged length of stay in the hospital ($>7$ days) for each patient. This task was run alongside MLM.
Pooling Strategies
C.4 Optimization Procedure
Each setting is introduced incrementally in the optimization procedure. We begin with the baseline model (BEHRT) and first introduce medication codes. If this modification enhances performance, it is incorporated into the model. Next, we test the second modification (Full ICD10/ATC Tree) on the updated model, which now includes the baseline and the improvement. This iterative process continues for each subsequent modification, with each new change tested on a model that incorporates all previous beneficial changes.
For evaluation, we use three generic downstream tasks for generalizable performance: one-year death prediction for patients older than 60 y.o., one-month pain treatment prediction, and one-week infection prediction; the corresponding definitions are given in Appendix S, consisting of diagnosis and medication codes as well as free text. Death encapsulates the final outcome and can signal the overall health status, encompassing various conditions and factors affecting patient well-being. Pain treatment is very frequent and covers a wide range of causes. It can be both mild and severe, indicative of many underlying conditions, and thus quite general. Infection points to a compromised immune system, which also tells us about the general condition of a patient.
Our three tasks constitute a generalizable evaluation by covering a broad range of health conditions. To compare the performance between models, we mainly considered the area under the receiver operating curve (AUROC) but also reported the area under the precision-recall curve (AUPRC) in Appendix R. For each downstream task, we perform five-fold cross-validation (CV) evaluating on a test set and compute the mean and standard deviation. The downstream task was a binary classification and was constructed as follows. For patients with positive outcomes, we used the outcome of interest as the index date. We then excluded all events within a specified prediction window leading up to the index date, with the window tailored to each downstream task. Disease definitions and prediction windows are given by expert clinicians, ensuring that the model sees no outcome-related information during prediction, preventing any information leakage. For patients with negative outcomes, we assign random censoring dates matching the distribution of the positive cases to avoid information leakage introduced by censoring. This makes the censoring identical for both cases and controls. To account for the imbalance in the dataset, we sampled with the inverse square root of the frequency of each class (Mahajan et al., 2018; Mikolov et al., 2013).
C.5 Assessing Model Generalizability
We also covered conditions with different associations to overall health and underlying factors. For instance, the strong correlation between stroke and overall health (Arboix, 2015) contrasts with the more complex interplay of genetic and environmental factors in schizophrenia (Stilo and Murray, 2019; Henriksen et al., 2017), which presents a subtler EHR data signature. We also covered conditions with a limited range of causes, such as anaphylaxis, which primarily involves drug-related causes, and those that can have a wide range of underlying conditions, such as sleep disorder. The conditions are listed in Table 5 together with prediction window and age groups, which were chosen based on relevance and considering the age distributions provided in Appendix P.
This structured approach to task selection is designed to demonstrate the generalizability of our model in identifying a wide range of health outcomes, underscoring its potential utility in diverse clinical scenarios without having to revisit the model choices for every new downstream task. We construct each downstream task as outlined in the previous subsection C.4, reusing the definitions from Appendix S, and perform hypothesis testing as described in Appendix H.
In addition, we performed experiments using out-of-time evaluation to mimic real-world scenarios more closely. Here, we look at temporally separated outcomes for train, validation, and test sets. A detailed description is provided in Appendix M. This method allows the model to predict the future and assess its response to temporal shifts.
C.6 Hyperparameters
To ensure a fair comparison, we standardized the hyperparameters across all experiments. We aimed to facilitate stable model convergence across all models while keeping manageable training durations. We utilized a base model architecture consisting of six layers and six attention heads with a hidden size of 192 and an intermediate size of 64. Sequences were truncated to 512 tokens, and a batch size of 256 for PT and 512 for FT was used. Learning rates were set to $10^{-3}$ for PT and $2\cdot 10^{-5}$ for FT using a linear warmup of five epochs. Models were trained for 50 epochs during optimization and 150 epochs for the generalization experiments with an early stopping of five epochs. A more detailed overview can be found in Appendix J.
Appendix D Cohort
@@
The data consists of EHRs from 1,825 million patients, prospectively recorded during the period from 2016 to the end of 2022 in the Capital Region of Denmark. We consider only patients with at least one code eligible, which results in 1,805 million patients. For FT, we deem at least one code to be a proper input for binary classification but use patients with $\geq 3$ codes for PT to ensure a proper input for the MLM objective task. An overview of the selection pipeline can be seen in Appendix N.
@@ D.1 Cohort Selection
To test generalizability across a wide range of tasks while managing computational demands, we maintained a manageable cohort size for the main experiments. However, we also performed ”maximizing performance” experiments, where we extended the cohorts, including PT patients in the FT set for select tasks. Details of these tasks and the expanded cohort sizes are provided in Appendix L.
The out-of-time evaluation, aimed at mimicking a real-world setting, was constructed by censoring training data on 01/01/2020. Without performing patient splits and keeping patients with at least one event we obtain 1,033,465 PT patients. The time separation is described in detail in Appendix M and patient numbers are provided in Table 3.
For model selection, we require each patient in the PT cohort to have at least three diagnoses to ensure a fair comparison between the diagnosis only and the diagnosis+medication models. For the final evaluation, we require at least three of any (diagnosis or medication) codes, to accurately reflect the increased cohort when including medication.
For each condition, we defined a relevant age and/or sex and a prediction window that we use for censoring, given in Table 1 for the optimization experiments and in Table 5 for the generalization assessment. The full age distributions for the cohorts before age group selection are given in Appendix P. We consider patients to be positive when having at least one of the codes given in Appendix S and negative otherwise. We address imbalance and potential information leakage using the methods described in subsection C.4.
D.2 Data Extraction
We extract all medical events, with an event consisting of three main variables: concept, timestamp, and admission ID. These events undergo minimal data processing through a simple three-stage process of replacing missing ICD10 codes with textual descriptions, attempting to impute missing values, only when confident, before excluding any events with missing info or ages outside a normal range. We leave a more detailed description in Appendix I.
D.3 Feature Choices
@@
Appendix E Results on Real Data
E.1 Evaluation Approach/Study Design